1. Which of the following statements are true and which are false? Give reasons for your

answer. (20)

i) The equation r = acosθ +bsinθ represents a circle.


ii) If 1,1/2,0 are direction ratios of a line, then the line makes an angle of 90◦ with the

x-axis, an angle of 60◦ with the y-axis, and is parallel to the z-axis.


iii) The intersection of a plane and a cone can be a pair of lines.


iv) If a cone has three mutually perpendicular generators then its reciprocal cone has

three mutually perpendicular tangent planes.


1
Expert's answer
2021-05-18T18:06:45-0400


i) TRUE


Explanation:


The equation r = acos "\theta" + bsin "\theta" represents a circle. 



"r=a\cos \theta+b\sin \theta"

Multiply both sides by r



"r^2=ar\cos\theta+br\sin\theta"

Convert to cartesian coordinates



"x^2+y^2=ax+by"

Complete the square



"x^2-ax+{a^2 \over 4}+y^2-by+{b^2 \over 4}={a^2 \over 4}+{b^2 \over 4}""(x-{a \over 2})^2+(y-{b \over 2})^2={a^2 \over 4}+{b^2 \over 4}"

Therefore, the equation r = r = acos "\theta" + bsin "\theta" represents a circle. 


ii) FALSE


Explanation:


 If 1,1/2,0 are direction ratios of a line, then the direction cosines will be as follows


l = "\cos\alpha" = 1 \ ( √ 12 + 0.52 + 02) = 2 / √5

m = "\cos\beta" = 0.5 \ ( √ 12 + 0.52 + 02) = 1 / √5

n = "\cos\gamma" = 0 \ ( √ 12 + 0.52 + 02) = 0


So the line makes 26.77"\degree" with the x - axis , 63.48"\degree" with the y - axis and 90"\degree" with the z - axis.


iii) TRUE


Explanation:


The intersection of a plane and a cone can be a pair of lines.

If you intersect a cone with a plane, the intersection will be one of the following: a parabola, a circle, an ellipse, a hyperbola, a pair of lines (the plane must lie along the axis of the cone), a single line (plane is tangent to cone), or a single unique point (the plane must be perpendicular to the axis, passing through the center).


iv) TRUE


EXPLANTION:


If a cone has three mutually perpendicular generators then its reciprocal cone has

three mutually perpendicular tangent planes and vice versa.